One of two complementary angles is 6 less than twice tthe other angle.Find the measure of the angles.
Answer by algebrahouse.com(1659)

"One of two complementary angles is 6 less than twice tthe other angle.Find the measure of the angles."

x = one angle
2x - 6 = other angle {six less than twice the other}

x + 2x - 6 = 90 {complimentary angles add up to 90 degrees}
3x - 6 = 90 {combined like terms}
3x = 96 {added 6 to both sides}
x = 32 {divided both sides by 3}
2x - 6 = 58 {substituted 32 for x, into 2x - 6}
www.algebrahouse.com